OpenStudy (anonymous):

use the power rule, the power of a product rule, and the power of a quotient rule to simplify the expression. (-4xz^5/y^2)^3 teach me please?

OpenStudy (anonymous):

(-4xz^5)^3 ( a )^n (a)^n (-------) first we use the rule (---) = ----- ( y^2 ) ( b ) (b)^n (-4xz^5)^3 ---------- now we use the rule (a^m)^n = a^mn ( y^2 )^3 -64x^3z^15 ----------- y^6 Hope it is clear.....☺

OpenStudy (anonymous):

The rules he is using are\[(\frac{a}{b})^{n} = \frac{a^{n}}{b^{n}}\]and\[(a^{m})^{n} = a^{mn}\]The other one you'll need is \[(ab)^{n} = a^{n}b^{n}\]

OpenStudy (anonymous):

Wow. Those exponents are completely illegible. If you increase your browser font size you can read them better.